TETSXPREME:

Jesse and Amir were assigned the same book to read. Jesse started reading on Saturday, and he is reading 30 pages a day. Amir didn't start until Sunday, but he is reading 35 pages a day. How many days will it take Amir to catch up to Jesse, and how many pages will they each have read? Answer the questions to solve this problem using a system of equations.

TETSXPREME:

Write an equation to represent the number of pages Amir has read. Use x to represent the number of days Amir has been reading and y to represent the number of pages he has read. Answer:35x=y

TETSXPREME:

Using x, the number of days Amir has been reading, write an expression to represent the number of days Jesse has been reading. Remember that Jesse started reading one day before Amir. Use this expression to write an equation for the number of pages Jesse has read. Let y represent the number of pages read. Answer:(35x-5x)+30=y

TETSXPREME:

Write the system of equations using your answers from questions 1 and 2.

YRJ8498:

O, amir needs to read 60 pages on sunday because jesse read 30 on saturday and read another 60 on sunday so amir needs to read 60 on sunday to catch up with jesse

YRJ8498:

Thats all im doing for you tet

TETSXPREME:

@tetsxpreme wrote:
Write the system of equations using your answers from questions 1 and 2.
i only needed help with this

YRJ8498:

@tetsxpreme wrote:
@tetsxpreme wrote:
Write the system of equations using your answers from questions 1 and 2.
i only needed help with this
YoU ShOuLd'Ve SaId ThAt

TETSXPREME:

i had the answer at the bottom

YRJ8498:

Are you in 7th grade?

TETSXPREME:

no 6th

YRJ8498:

Bruh ask your teacher it is kind confusing

YRJ8498:

tell me what she tells you

TETSXPREME:

its like alot of numbers 233xw+3 2345_2x+ =22

TETSXPREME:

its something like that but idrk

TETSXPREME:

thats how its supposed to look like

jhonyy9:

- my idea Jesse start on Saturday with 30 pages a day so in 5 days (for example ) read 150 pages Amir start on Sunday with 35 pages a day so on this (5-1=4) days will read 4*35=140 pages so in 10 days Jesse read 10*30 = 300 pages what mean for Amir 10-1=9 days = 9*35 = 315 pages - any idea now about these above ?

TETSXPREME:

is that for the 3rd question?

ISayYuh:

Did you find the answer? he will have to read more than thirty pages on sunday because jesse already read 60 pages